OpenStudy (anonymous):

the shorter leg of a 30-60-90 triangle is 6 what is the length of the hypotenuse? @DebbieG

OpenStudy (yttrium):

 Since it is stated that the shorter leg is 6, it means that it is found at the opposite of the shorter angle, therefore it must be placed at the opposite of 30 degrees, With that you can use the pythgagorean identies. Can you do it yourself? Or you need further help?

OpenStudy (anonymous):

help

OpenStudy (yttrium):

Wait, did you get what I want to explain? Before we continue.

OpenStudy (anonymous):

sorta

OpenStudy (yttrium):

We all know that sin = opp/hyp, right? So we can use this find the length of the hypotenuse. That is, sin30 = 6/hyp Therefore, hyp = 6/sin30 hyp = 12

OpenStudy (yttrium):

I mean the 12 in the drawing is 6 huh. ^^

OpenStudy (anonymous):

yes i put it in the calcuator and got 12

OpenStudy (yttrium):

Therefore, what is the answer? :)

OpenStudy (anonymous):

12

OpenStudy (anonymous):

12

OpenStudy (yttrium):

Right!

OpenStudy (anonymous):

i have a new question i just posted i am kindda lost i got to the second step and got stuck